# Zero-Shot Clinical Trial Patient Matching with LLMs

### 08 Apr 2024 v2

@@ Abstract
Matching patients to clinical trials is a key unsolved challenge in bringing new drugs to market. Today, identifying patients who meet a trial’s eligibility criteria is highly manual, taking up to 1 hour per patient. Automated screening is challenging, however, as it requires understanding unstructured clinical text. Large language models (LLMs) offer a promising solution. In this work, we explore their application to trial matching. First, we design an LLM-based system which, given a patient’s medical history as unstructured clinical text, evaluates whether that patient meets a set of inclusion criteria (also specified as free text). Our zero-shot system achieves state-of-the-art scores on the n2c2 2018 cohort selection benchmark. Second, we improve the data and cost efficiency of our method by identifying a prompting strategy which matches patients an order of magnitude faster and more cheaply than the status quo, and develop a two-stage retrieval pipeline that reduces the number of tokens processed by up to a third while retaining high performance. Third, we evaluate the interpretability of our system by having clinicians evaluate the natural language justifications generated by the LLM for each eligibility decision, and show that it can output coherent explanations for 97% of its correct decisions and 75% of its incorrect ones. Our results establish the feasibility of using LLMs to accelerate clinical trial operations.
@@ 1 Introduction
In this paper, we aim to build upon recent work exploring the application of LLMs to patient matching (Jin et al., 2023; Wong et al., 2023; Nievas et al., 2023). In contrast to this prior work, our primary focus is on the efficiency of LLM-based approaches – namely, their cost in terms of dollars, money, and time. These considerations are essential if such systems aim to be deployed at scale given the limited resources of most health systems.

### 10 Apr 2024 v3

@@ Abstract
Background: Matching patients to clinical trials is a key unsolved challenge in bringing new drugs to market. Today, identifying patients who meet a trial’s eligibility criteria is highly manual, taking up to 1 hour per patient. Automated screening is challenging, however, as it requires understanding unstructured clinical text.
Methods: We design a zero-shot LLM-based system which, given a patient’s medical history as unstructured clinical text, evaluates whether that patient meets a set of trial inclusion criteria (also specified as free text). We investigate different prompting strategies and design a novel two-stage retrieval pipeline to reduce the number of tokens processed by up to a third while retaining high performance.
Results: We achieve state-of-the-art performance on the n2c2 2018 cohort selection challenge, the largest clinical trial patient matching public benchmark. Second, we show that our system can improve the data and cost efficiency of matching patients an order of magnitude faster and more cheaply than the status quo. Third, we measure the interpretability of our system by having clinicians evaluate the natural language justifications generated for each eligibility decision, and show that it can output coherent explanations for 97% of its correct decisions and even 75% of its incorrect ones.
Conclusion: Our results establish the feasibility of using LLMs to accelerate clinical trial operations. Our zero-shot retrieval architecture enables scaling to arbitrary trials and patient record length with minimal reconfiguration.
@@ 1 Introduction
In this paper, we aim to build upon recent work exploring the application of LLMs to patient matching (Jin et al., 2023; Wong et al., 2023; Nievas et al., 2023). In contrast to this prior work, our primary focus is on the efficiency of LLM-based approaches – namely, their cost in terms of data, money, and time. These considerations are essential if such systems aim to be deployed at scale given the limited resources of most health systems.